A rectangular lawn has the following description:
Length; 60 yards.
Width; 40 yds.
A 6 ft. wide walkway surrounding lawn, not included in overall dimensions.
Paths of same width are through lawn, connecting on opposite sides.
Determine area in sq. yds. 1 of the 4 plots enclosed by the paths.

Not sure how to solve.

Answer by ikleyn(52905)

.

From the area of 40*60 = 2400 sq.yards subtract the area

    60*6 = 360 sq.yards of one strip 

and subtract the area

    40*6 = 240 sq.yards of the other strip.



You will get  then  2400 - 360 - 240 = 2400 - 600 = 1800 sq.yards.



To this, add the area  6*6 = 36 sq.yards of the intersection square,
since you subtracted it twice.


By doing this way, you will get the final 



ANSWER.  1836 sq.yards.


Solved, answered, explained and completed.